OpenStudy (anonymous):

I am having some trouble figuring out how to approach the following problem: A 75.0-mL volume of 0.200 M NH3 (Kb=1.8×10−5) is titrated with 0.500 M HNO3. Calculate the pH after the addition of 17.0mL of HNO3.

OpenStudy (kira_yamato):

We look at the end of the day: n(HNO3) added = 0.500*17.0/1000 = 0.00850 mol n(NH3) = 0.200*75.0/1000 - 0.00850 = 0.00650 mol [NH3] left = 0.00650*1000/(17.0+75.0) = 0.070652 M [OH-] = Kb * [NH3] = 0.070652*1.8*10^(-5) = 1.27174 x 10^(-6) pOH = -log[OH-] ≈ 5.8956 pH = 14 - pOH ≈ 8.10